APACHE POINT OBSERVATORY SDSS 2.5M OBSERVING LOG Tuesday February 9, 2010 (MJD 55237) ---=== OBSERVING TEAM ===--- Steph Snedden Alaina Shelden Kaike Pan (afternoon) ---=== OBSERVING PLAN ===--- BOSS Science! ---=== OBSERVING SUMMARY ===--- Night lost to clouds and humidity. The spectro crate looks like it was shut down...? ---=== OBSERVING LOG ===--- Afternoon: ---------- All systems, actors, telescope were checked. No problem to report. Paul is working on guider, and running trunk version guiderActor. Boss collimation was checked with cartridge 17. Hight humidity prevented afternoon ventilation. Night: ------ 00:53Z Humidity has dropped, but we still have alot fo clouds. 01:03Z Started ventilation, on the chance that the clouds might clear up. 01:57Z boss exposure dark itime=20 02:03Z sop doCalibs nflat=1 02:07Z sop doCalibs narc=1 02:09Z Gary did some work today on sos, now have a number for plate 3769. Updated database. Problems from last night fixed. 02:27Z Stopped ventilation, humidity high again. 11:50Z endNight finished manually, and there was much rejoicing. Yay. ---=== IMAGING RUN SUMMARY ===--- Run Time Stripe Lambda Last Flavor Comments Start End Begin End Frame ------------------------------------------------------------------------- ---=== IMAGING RUN DETAILS ===--- ---=== SKIPPY RESULTS ===--- Run Frame nFrames stars muErr muRms nuErr nuRms rot az el --------------------------------------------------------------------------- ---=== LTMATCH RESULTS ===--- Run Field nFields alt az nGood rowMean rowSig colMean colSig rot ------------------------------------------------------------------------ ---=== SPECTROSCOPY DATA SUMMARY ===--- Summary Checked (y/n): QA Procedures Done (y/n): UT Exp Time flavor comment (S/N)^2 totals ========================================== b1 r1 b2 r2 ---=== TELESCOPE OFFSETS AND SCALE I ===--- Time Instrument Az Alt Rot Scale pos offset pos offset pos offset ------------------------------------------------------------------------------ ---=== TELESCOPE OFFSETS AND SCALE II ===--- ---=== FOCUS LOG ===--- setmir piston Temp Wind Time Inst scale M1 M2 Foc Az Alt (C) MPH Dir filt fwhm ------------------------------------------------------------------------------ ---=== WEATHER LOG ===--- Wind Time Temp F Dewp F MPH Direction Dust DIMM Sky 22:27Z 27 20 6 242 (WSW) 501 - clouds, humid 22:59Z 26 21 9 224 (SW) 510 - 23:33Z 24 20 8 187 (S) 609 - 00:55Z 24 10 14 197 (SSW) 426 - 01:28Z 25 13 14 187 (S) 453 - 02:01Z 25 18 15 180 (S) 567 - 02:34Z 25 21 8 181 (S) 642 - 03:07Z 25 22 14 166 (SSE) 718 - 03:40Z 25 22 16 180 (S) 608 - 04:13Z 25 22 4 192 (SSW) 652 - 04:46Z 25 21 15 191 (S) 558 - 05:18Z 25 22 11 185 (S) 593 - 05:51Z 25 22 7 183 (S) 630 - 06:24Z 25 22 8 188 (S) 594 - 06:57Z 26 23 13 182 (S) 643 - 07:30Z 26 23 9 186 (S) 623 - 08:03Z 26 24 15 168 (SSE) 664 - 08:36Z 27 23 16 186 (S) 627 - 09:09Z 27 24 12 185 (S) 649 - 09:42Z 27 24 16 189 (S) 684 - 10:15Z 26 24 15 209 (SSW) 817 - 10:48Z 26 24 9 180 (S) 740 - 11:21Z 25 23 9 189 (S) 752 - 11:54Z 25 24 9 192 (SSW) 682 - clouds humid ---=== TELESCOPE STATUS ===--- 01:03Z Started ventilation 02:27Z Stopped ventilation. No interlocks bypassed. Status at: 11:53Z Telescope stowed at: 30,121,00 mount Instrument mounted: Cartridge 17 Counterweights at: 240 lbs LN2 autofill systems: Connected and turned on. 180L LN2 dewar scales: Spectro 155.1 lbs, 19.3 psi Imager NA lbs, NA psi ---=== SOFTWARE USED ===--- IOP/SOP: v4_40_0 Watcher: v2_36_0 MCP: v6_6_3 TPM: tpm_v3_5_0 AstroDa: v15_13_1 TCC: TCC 2.11.0 2009-06-24 sopActor: v1_0_4 guiderActor: new truck version gcameraICC: v1_0_5 alertsActor: v1_0_3 platedbActor: v1_0_2 platedb v1_0_3 sopActor: /home/sdss3/products/Linux64/sopActor/v1_0_4 guiderActor: /home/sdss3/products/Linux64/guiderActor/v1_0_16 gcameraICC: /home/sdss3/products/Linux64/gcameraICC/v1_0_5 alertsActor: /home/sdss3/products/Linux64/alertsActor/v1_0_3 platedbActor: /home/sdss3/products/Linux64/platedbActor/v1_0_2 ---=== MIRROR NUMBERS ===--- PRIMARY: -------- Scale: 1.000000 MIGS TONIGHT NOMINAL Axial A -0.7050 -0.6940 Axial B -0.0280 -0.0410 Axial C -0.3790 -0.3710 Trans D 11.6060 11.6170 Lateral E 10.3759 10.3505 Lateral F 11.9634 11.9253 GALILS Commanded: 9900. 400. 5850. 1200. 2950. 2750. Actual: 9917. 401. 5832. 1205. 2954. 2757. SETMIR VALUES PriDesOrient: 0.00 -3.30 5.10 278.08 178.06 PriOrient: 0.00 -3.04 5.26 277.70 178.07 SECONDARY: ---------- Focus: 0.00 Air Temp.: -2.7 Alt.: 29.999997 MIGS TONIGHT NOMINAL Axial A 18.6170 18.6150 Axial B 15.6300 15.6120 Axial C 13.8290 13.8170 Trans D -11.7770 -11.7780 GALILS Commanded: 1600329. 1564625. 1585898. -7600. -12400. Actual: 1597721. 1566188. 1588095. -7597. -12380. SETMIR VALUES SecDesOrient: 1257.00 -4.00 -9.00 10.00 234.42 SecOrient: 1256.69 -3.31 -10.43 9.25 234.24 ---=== PROBLEMS IN DETAIL ===---